OpenStudy (anonymous):

Simplify and find the domain of : sin(2(sin(2x))), I have 2sin(sin(2x))cos(sin(2x))

OpenStudy (anonymous):

and for the domain I have \[\mathbb{R} \] except for \[(3\Pi/4)+(\Pi n)\]

OpenStudy (anonymous):

the domain is all real numbers

OpenStudy (anonymous):

i have no idea what "simplify" means in this context, but you can take the sine of anything. there are no restrictions

OpenStudy (anonymous):

but what if 2x = 270º, sin(270) = -1, ahhhh; BUT we dont care, lol, I was mixed up with another number

OpenStudy (anonymous):

good. makes no difference at all

OpenStudy (anonymous):

oh, ok, well. 1 sec. Ill post the question as it is in my text book. just a sec

OpenStudy (anonymous):

it is g(g(x)) and g(x) = sin(2x). so what i did was sin(2(sin(2x)))

OpenStudy (anonymous):

you can rewrite this beast as \[2 \sin(\sin(2 x)) \cos(\sin(2 x))\] if you like not sure why that is simpler though

OpenStudy (anonymous):

yes you are correct.

OpenStudy (anonymous):

ok, I see. Well. Thanks very much, my question is answered.